Question 32 1 pts
An ANOVA is used to evaluate the mean differences among three treatment conditions with a sample of n = 12 participants in each treatment. For this study, what is dfwithin?
11
Answer #1

32) 33

number of treatments (K)-3

Number of sample participants (n) -12

Number of scores in the total research study (N) -36 (n*k = 3 * 12= 36)

df(within) is N - k , which is 33 (36-3).